If Tom bought 11 cans of cooking oil and each can had a radius of 5.4 inches and was 11 inches tall, how many total cubic inches of cooking oil would Tom have?
Answer by checkley75(3666)

VOLUME OF THE CAN IS V=PIR^2*H
V=3.14(5.4^2)11
V=3.14*29.16*11
V=1007.19 IN^3 IS THE VOLUME OF A CAN.
11*1007.19=11079.09 IN^3 IS THE TOTAL VOLUME OF COOKING OIL.